# Debating Truth: Debate-driven Claim Verification with Multiple Large Language Model Agents

### v2

@@ Abstract.
Claim verification is essential for digital literacy, yet state-of-the-art single-agent methods often struggle with complex claims that require nuanced analysis of multifaceted online evidence. Inspired by real-world human fact-checking practices, we propose DebateCV, the first debate-driven claim verification framework powered by multiple LLM agents. In DebateCV, two Debaters argue opposing stances over multiple rounds to surface subtle errors in single-agent assessments. A decisive Moderator is then required to weigh the evidential strength of conflicting arguments to deliver an accurate verdict. Yet zero-shot agents struggle to adjudicate multi-round debates for verifying complex claims, often defaulting to neutral judgements, and no datasets exist for training agents for this role. To bridge this gap, we propose Debate-SFT, a post-training framework that leverages synthetic data to enhance agents’ ability to effectively adjudicate debates for claim verification. Results show that our methods surpass state-of-the-art non-debate approaches in both accuracy (across various evidence conditions) and justification quality, which strengthens societal resilience against misinformation and contributes to a more trustworthy online information ecosystem.
Responsible Web AI, Web Technologies for Combating Misinformation, Debate-driven Claim Verification, Automated Fact-Checking
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of The Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesisbn: 978-1-4503-XXXX-X/2026/04ccs: Information systems Data miningccs: Computing methodologies Natural language processingccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
In the modern digital landscape where misinformation disseminates widely and rapidly online (factcheck_definition; miyazaki2023fake; zhang-2025-generation; active; RumorGraphXplainer), (automated) claim verification has become a cornerstone web technology for assessing the veracity of (check-worthy) claims and promoting digital literacy (factcheck_definition; guo2022survey). [^1] Check-worthy claims are statements that can impact public opinion or behaviour (annotation_schema). State-of-the-art (SOTA) claim verification approaches typically use the retrieval-augmented generation (RAG) framework (averitec_challenge), wherein a large language model (LLM) agent receives a prompt that integrates the claim and relevant evidence retrieved from the web, and subsequently generates a verdict along with justifications that elucidate the rationale behind its veracity assessment.
@@
However, the claim verification capabilities of individual agents are insufficient for complex claims that involve multifaceted evidence (conflicts; mohsin2025fundamental). Take HerO, the SOTA method using a single agent with supervised fine-tuning (SFT) and Chain-of-Thought (CoT) prompting, as an example. It incorrectly refutes the claim “52% of Nigeria’s current population lives in urban areas” by relying on outdated evidence indicating a 36% urban population in 2008. Although more recent evidence supporting a 52% urbanization rate was successfully retrieved, the single agent in this approach completely overlooked this critical evidence, resulting in this error. [^2] This error traces back to Claim 218 in hero’s official results.
@@
To mitigate errors in individual claim verifications, professional fact-checking agencies in real-world practice commonly adopt a debate-driven claim verification methodology. For instance, factcheck_ecology describes how PolitiFact, a Pulitzer Prize-winning fact-checking agency, employs “star chamber” sessions, i.e., debates among fact-checkers, to verify claims. [^3] Please see examples of real-world claim verification debates in (factcheck_ecology), pp. 204–208. In these sessions, a panel of fact-checkers critically reviews each other’s veracity assessments to identify potential flaws in evidence analysis, e.g., the oversight of the latest urbanization statistics in the above example. Such an adversarial verification process compels the fact-checkers to iteratively refine their assessments, thereby enhancing the depth and rigour of evidence analysis and fostering holistic claim evaluation (factcheck_ecology).
@@
Inspired by these real-world debate-driven practices, this work introduces DebateCV, the first debate-driven claim verification framework based on multiple LLM agents. Specifically, DebateCV employs two role-playing Debater agents with opposing stances: one affirming and the other refuting the claim’s veracity. These Debaters iteratively present arguments, contest each other’s positions, and defend their own viewpoints by leveraging collected evidence over multiple rounds. In each round, a third Moderator agent evaluates the presented arguments and decides whether further debate is necessary. If the debate concludes, the Moderator is tasked with assessing the evidential strength of both sides across all rounds and synthesising an accurate, well-justified verdict. However, our preliminary experiments (see Sec. 5.3) show that, while zero-shot LLM agents perform effectively as Debaters, they struggle in the Moderator role. For complex claims requiring multiple rounds of debate, they often fail to identify a clear winner, even when sufficient evidence supports a decisive true/false verdict, and instead frequently default to neutral outcomes, such as stating that the evidence is insufficient. This limitation calls for a dataset containing complete claim verification debate transcripts paired with expert verdicts and justifications grounded in the debates to enable SFT of the Moderator agent. Unfortunately, no such dataset exists, and manual construction is prohibitively expensive. In response, we introduce Debate-SFT, the first post-training framework specifically designed to fine-tune the Moderator for debate-driven claim verification using synthetic debate data. As illustrated in Fig. 1, our framework begins by constructing the SynDeC (Synthetic Debate for Claim verification) dataset, which is generated by executing zero-shot DebateCV to synthesise multi-round debate transcripts for various claims. We then identify and correct discrepancies between the zero-shot Moderator’s outputs and ground-truth labels to ensure a clean training signal. The Moderator is subsequently post-trained on the curated SynDeC dataset to enhance nuanced argument evaluation.
@@
With the enhanced Moderator, DebateCV leverages a simulated “star chamber” among LLM agents to rigorously analyse evidence from multiple perspectives. Experiments show that DebateCV outperforms existing SOTAs by 2.6–5.8% in accuracy across diverse evidence quality settings. Case studies illustrate that DebateCV’s adversarial debate mechanism effectively corrects typical single-agent errors, such as claim/evidence misinterpretation, overlooked evidence, and overreliance on speculation. Human expert evaluations further confirm that DebateCV provides higher-quality justifications than previous methods, beyond improvements in accuracy.
@@ 2. Related Work
Automated fact-checking systems (infact; hero; CTU_at_averitec; competing_wisdom; guo2022survey) typically operate in two stages: first, they retrieve relevant evidence (evidence retrieval stage); then, they use a single LLM agent to predict a verdict, such as Supported, Refuted, Not Enough Evidence, or Conflicting Evidence/Cherry-picking, along with a justification (averitec_challenge; hero) (claim verification stage). [^4] Appendix A provides the definition of the four verdict categories.
However, as shown in Sec. 1, existing single-agent systems are vulnerable to multifaceted and even misleading evidence retrieved from the open web. This vulnerability can arise from a fundamental mismatch: LLMs are pretrained to model linguistic patterns by autoregressively predicting the next token, while claim verification demands comprehensive analysis of collected evidence (mohsin2025fundamental). Ensembling multiple agents, e.g., via majority voting (debate_or_vote; fact_audit) or collaborative frameworks (metagpt; camel), also lacks rigorous adversarial scrutiny and may amplify individual errors instead of correcting them. Inspired by real-world fact-checking practices (factcheck_ecology), we propose the first debate-driven claim verification framework, DebateCV, to expose and resolve individual assessment errors.
Close to our work, a few recent studies (competing_wisdom; yue2024retrieval) have explored the value of integrating evidence or arguments from opposing perspectives. competing_wisdom retrieve and present separate sets of supporting and opposing evidence to a single LLM for verdict prediction, which focuses on improving evidence retrieval. Instead, we target the subsequent claim verification stage by enabling comprehensive analysis of retrieved evidence. yue2024retrieval generate static supporting and refuting arguments independently, but their approach lacks an interactive debate mechanism to surface flaws or resolve contradictions, a mechanism that DebateCV mimics from human fact-checking. Another concurrent work (han2025beyond) investigates evidence-based multi-agent debate for misinformation detection, but it is limited to binary (true/false) classification. This oversimplifies real-world scenarios where claims may lack sufficient evidence (Not Enough Evidence) or involve ambiguity and source disagreement (Conflicting Evidence/Cherry-picking). Also, they rely solely on prompting zero-shot LLMs, which, as shown in Sec. 5.3, exhibit conformity bias and frequently default to neutral verdicts in realistic non-binary settings. In contrast, we introduce Debate-SFT, a novel post-training framework that substantially improves LLMs’ ability to adjudicate complex debates.
@@
Recent studies have shown that multi-agent interactions empower diverse applications, such as AI societies (camel) and software engineering (autogen; metagpt), with multi-agent debate-based frameworks often outperforming other forms in complex tasks, such as machine translation (mad), hallucination mitigation (debate_improve_factuality; debaterag), and stance detection (lan2024stance). However, prior approaches to misinformation detection (debate_for_fake; debate_to_detect; liu2025truth) mainly debate linguistic patterns of fake news without retrieving external evidence. These patterns typically reflect platform-specific (www_mcfend) or time-constrained features (yang2024search), limiting their real-world applicability. In contrast, this work targets the realistic claim verification task that analyses consistencies and contradictions between retrieved evidence and claims (guo2022survey), and is the first to propose a debate-driven framework for this task.
Beyond downstream applications, three prior studies have explored leveraging debate mechanisms for post-training LLMs. debategpt employ debates to refine training corpus annotations for general language modelling. debat_for_diverse use debates to generate convincing arguments for fine-tuning LLM persuasiveness. acc_dabte use debate games between two LLM teams to improve their team-working abilities. Different from these works, we propose Debate-SFT, the first post-training framework specifically designed for enhancing LLMs’ ability to distil multi-round debates into accurate verdicts. We note that, though tailored for claim verification, Debate-SFT, with minimal modifications, can be adapted to other contexts using multi-agent debate.
@@ 3. DebateCV: Debate-driven Claim Verification
Debate-driven claim verification is inspired by human fact-checking, where multi-round debates enable thorough evidence analysis. We formalize this task as determining the veracity of a textual claim $c$ using an evidence set $\mathcal{E}=\{e_{1},\ldots,e_{k}\}$ by simulating a debate between specialised LLM agents. The outcome is a verdict $\hat{y}$ and a corresponding justification $\hat{j}$, both grounded in the arguments exchanged during the debate. In this section, we present DebateCV and describe how LLM agents are coordinated within the framework.
@@ 3.1. Pre-debate Configuration
Similar to debates conducted by humans, debates conducted by LLM agents must be carefully configured through role assignment. DebateCV employs three distinct LLM agents with complementary responsibilities: two Debaters, namely an Affirmative Debater $D^{+}$ and a Negative Debater $D^{-}$, and one Moderator $M$. Each agent is initialized with tailored system prompts that define and guide their behaviour throughout the debate. Their roles are outlined as follows.
@@ 3.2. In-Debate Process
At the end of each round, the Moderator ($M$) synthesises the key arguments and evidence from both sides into a summary of all exchanged arguments ($S_{t}$). The Moderator then assesses whether the debate has converged, e.g., it deems that $S_{t}$ and $S_{t-1}$ are highly similar, indicating no new insights are being introduced. The debate terminates early if $M$ detects convergence; otherwise, it proceeds until the maximum number of rounds ($T_{\text{max}}$) is reached. Upon conclusion, $M$ determines the final verdict ($\hat{y}$) by weighing the arguments from both sides and provides a justification ($\hat{j}$) based on the debate.
Algorithm 1 presents a high-level overview of DebateCV. The specific prompts used for each agent are provided in Appendix B.
Algorithm 1 Pseudocode Summary of DebateCV.
1:Claim $c$, evidence set $\mathcal{E}=\{e_{1},\dots,e_{k}\}$, round limit $T_{\text{max}}$
2:Verdict $\hat{y}$, justification $\hat{j}$
3:Initialize $^{+}, DD^{-}$, $M$ with their roles, $c$ and $\mathcal{E}$
[…]
5:while $t\leq T_{\text{max}}$ do
6:  if $t=1$ then
7:   $D^{+}$ generates $A^{+}_{1}$ supporting $c$ using $\mathcal{E}$
8:  else
9:   $D^{+}$ generates $A^{+}_{t}$ defending against $A^{-}_{t-1}$ using $\mathcal{E}$
10:  end if
11:  $D^{-}$ generates rebuttal $A^{-}_{t}$ challenging $A^{+}_{t}$ using $\mathcal{E}$
12:  $M$ synthesises a summary $S_{t}$ for all arguments so far
13:  if $t>1$ and $M$ detects convergence (e.g., $S_{t}\approx S_{t-1}$) then
14:   break
15:  end if
16:  $t\leftarrow t+1$
17:end while
18:$M$ outputs final verdict $\hat{y}$ and justification $\hat{j}$
19:return $\hat{y},\hat{j}$
@@ 4. Debate-SFT: Post-training the Moderator
The DebateCV framework, introduced in Sec. 3, provides a debate-driven approach to claim verification. In our preliminary experiments, we find that zero-shot Debater agents consistently produce well-reasoned arguments from their assigned perspectives (see Sec. 5.4 for human evaluation results regarding debate quality). However, the zero-shot Moderator often struggles with complex claims that require multiple debate rounds and tends to exhibit conformity bias (conform1; conform2). Specifically, the Moderator frequently defaults to neutral verdicts such as Not Enough Evidence or Conflicting Evidence/Cherry-picking, even when sufficient evidence exists to support a definitive judgement (see Sec. 5.3 for experimental results).
To enhance the LLM agents’ ability to adjudicate debates, we aim to employ supervised fine-tuning (SFT), a well-established approach for adapting LLMs to downstream tasks (post-train_survey). However, SFT requires annotated debate-driven claim verification datasets, which contain complete multi-round debate transcripts paired with final verdicts and justifications grounded in the debates. Such datasets are both scarce and expensive to curate. To address this challenge, we introduce Debate-SFT, a novel post-training framework that transforms existing non-debate claim verification datasets into synthetic debate-driven equivalents, enabling effective fine-tuning of the Moderator. This section outlines the three stages of Debate-SFT: (1) debate data synthesis, (2) error correction, and (3) post-training.
@@
Given a non-debate claim verification dataset with $n$ claims $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, their ground-truth verdicts $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$, and golden evidence sets $\mathcal{E}=\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{n}\}$ annotated by humans, we leverage the zero-shot DebateCV framework (Algorithm 1) to generate synthetic debate recordings $\mathcal{D}=\{d_{1},d_{2},\dots,d_{n}\}$. Each recording $d_{i}$ is a multi-round transcript of the arguments and summaries produced during the debate over claim $c_{i}$: $d_{i}=\bigl\{A^{+}_{1},A^{-}_{1},\dots,A^{+}_{t_{i}},A^{-}_{t_{i}}\bigr\},$ where $A^{+}_{t_{i}}$ and $A^{-}_{t_{i}}$ are the affirmative and negative arguments in round $t_{i}$}. [^5] Our human evaluation in Sec. 5.4 shows that the generated debates feature well-reasoned, evidence-based arguments and consistently uphold opposing stances. During this process, the zero-shot Moderator predicts verdicts $\mathcal{\hat{Y}}=\{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{n}\}$ with justifications $\mathcal{\hat{J}}=\{\hat{j}_{1},\hat{j}_{2},\dots,\hat{j}_{n}\}$ based on $\mathcal{D}$.
@@
Since the zero-shot Moderator may issue erroneous verdicts even when the debaters’ arguments are established in gold-standard evidence, we use human-annotated ground-truth verdicts to ensure data validity. Specifically, when the predicted verdict $\hat{y}_{i}$ differs from the ground-truth $y_{i}$, we introduce an LLM-based Corrector to generate a justification that aligns with the human annotation. For each claim $c_{i}$ where $\hat{y}_{i}\neq y_{i}$, the Corrector takes the full debate recording $d_{i}$ and the human-annotated ground-truth verdict $y_{i}$ as input to generate a corrected justification $j^{\prime}_{i}$ that explains how $y_{i}$ can be derived from the arguments and evidence in $d_{i}$. [^6] Our human evaluation in Sec. 5.4 shows that these Corrector-generated justifications effectively reflect the debated content and align with human judgements on the verdict.
@@
where $\mathcal{L}_{\text{SFT}}$ is the standard SFT loss (sft) and $\theta$ represents the zero-shot Moderator parameters. This three-stage Debate-SFT framework results in an enhanced Moderator that can better evaluate debate dynamics and render more accurate claim verification judgements.
@@
To instantiate the Debate-SFT framework, we apply it to the training split of AVeriTeC (averitec) (3,068 claims) to create the first synthetic dataset designed for debate-driven claim verification, SynDeC. [^7] Our method can be applied to any non-debate claim verification dataset with human-annotated verdicts and evidence. We select AVeriTeC for constructing SynDeC because it is the only real-world claim verification benchmark that rigorously addresses evidence leakage and insufficiency issues prevalent in prior datasets (defame; fact2fiction). Specifically, we use the zero-shot DebateCV framework to synthesise full debate recordings $d_{i}$ for all 3,068 claims. Among these, 916 claims are incorrectly predicted and undergo error correction, where a Corrector based on GPT-4o generates a revised justification $j^{\prime}_{i}$ aligned with the ground-truth verdict $y_{i}$. Each sample in the SynDeC dataset contains the following elements: (i) claim $c_{i}$, (ii) evidence set $\mathcal{E}_{i}$, (iii) debate recording $d_{i}$, (iv) ground-truth verdict $y_{i}$, (v) predicted verdict $\hat{y}_{i}$, (vi) predicted justification $\hat{j}_{i}$, and, when $\hat{y}_{i}\neq y_{i}$, (vii) the corrected justification $j^{\prime}_{i}$.

### v3

@@ Abstract.
State-of-the-art single-agent claim verification methods struggle with complex claims that require nuanced analysis of multifaceted evidence. Inspired by real-world professional fact-checkers, we propose DebateCV, the first debate-driven claim verification framework powered by multiple LLM agents. In DebateCV, two Debaters argue opposing stances to surface subtle errors in single-agent assessments. A decisive Moderator is then required to weigh the evidential strength of conflicting arguments to deliver an accurate verdict. Yet, zero-shot Moderators are biased toward neutral judgments, and no datasets exist for training them. To bridge this gap, we propose Debate-SFT, a post-training framework that leverages synthetic data to enhance agents’ ability to effectively adjudicate debates for claim verification. Results show that our methods surpass state-of-the-art non-debate approaches in both accuracy (across various evidence conditions) and justification quality.
Debate-driven Claim Verification, Automated Fact-Checking
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792993isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Data miningccs: Computing methodologies Natural language processingccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
In the modern digital landscape where misinformation disseminates widely and rapidly online (Vlachos and Riedel, 2014), (automated) claim verification has become a cornerstone web technology for assessing the veracity of (check-worthy) claims (Vlachos and Riedel, 2014; Guo et al., 2022; Konstantinovskiy et al., 2021). State-of-the-art (SOTA) claim verification approaches typically use retrieval-augmented generation (RAG) (Schlichtkrull et al., 2024a), wherein a large language model (LLM) agent receives a claim and relevant evidence retrieved from the web, and generates a verdict with explanatory justifications.
@@
However, the claim verification capabilities of individual agents are insufficient for complex claims that involve multifaceted evidence (Ge et al., 2025; Mohsin et al., 2025). Take HerO, the SOTA method using a single agent with supervised fine-tuning (SFT) and Chain-of-Thought (CoT) prompting, as an example. It incorrectly refutes the claim “52% of Nigeria’s current population lives in urban areas” by relying on outdated evidence indicating a 36% urban population in 2008. Although more recent evidence supporting a 52% urbanization rate was successfully retrieved, the single agent in this approach completely overlooked this critical evidence, resulting in this error. [^1] This error traces back to Claim 218 in Yoon et al. (2024)’s official results.
@@
To mitigate errors in individual claim verifications, professional fact-checking agencies in real-world practice commonly adopt a debate-driven claim verification methodology. For instance, Graves (2013) describes how PolitiFact, a Pulitzer Prize-winning fact-checking agency, employs “star chamber” sessions, i.e., debates among fact-checkers, to verify claims. [^2] Please see examples of real-world claim verification debates in (Graves, 2013), pp. 204–208. In these sessions, a panel of fact-checkers critically reviews each other’s veracity assessments to identify potential flaws in evidence analysis, e.g., the oversight of the latest urbanization statistics in the above example. Such an adversarial verification process compels the fact-checkers to iteratively refine their assessments, thereby enhancing the depth and rigour of evidence analysis and fostering holistic claim evaluation (Graves, 2013).
@@
Inspired by these real-world debate-driven practices, this work introduces DebateCV, the first debate-driven claim verification framework based on multiple LLM agents. [^3] The initial version of this work was completed in February 2025 (He et al., 2025). Specifically, DebateCV employs two role-playing Debater agents with opposing stances: one affirming and the other refuting the claim’s veracity. These Debaters iteratively present arguments, contest each other’s positions, and defend their own viewpoints by leveraging collected evidence over multiple rounds. In each round, a third Moderator agent evaluates the presented arguments and decides whether further debate is necessary. If the debate concludes, the Moderator is tasked with assessing the evidential strength of both sides across all rounds and synthesising an accurate, well-justified verdict.
However, our preliminary experiments (see Sec. 5.3) show that, while zero-shot LLM agents perform effectively as Debaters, they struggle in the Moderator role. For complex claims requiring multiple rounds of debate, they often fail to identify a clear winner, even when sufficient evidence supports a decisive true/false verdict, and instead frequently default to neutral outcomes, such as stating that the evidence is insufficient. This limitation calls for a dataset containing complete claim verification debate transcripts paired with expert verdicts and justifications grounded in the debates to enable SFT of the Moderator agent. Unfortunately, no such dataset exists, and manual construction is prohibitively expensive. In response, we introduce Debate-SFT, the first post-training framework specifically designed to fine-tune the Moderator for debate-driven claim verification using synthetic debate data. As illustrated in Fig. 1, our framework begins by constructing the SynDeC (Synthetic Debate for Claim verification) dataset, which is generated by executing zero-shot DebateCV to synthesise multi-round debate transcripts for various claims. We then identify and correct discrepancies between the zero-shot Moderator’s outputs and ground-truth labels to ensure a clean training signal. The Moderator is subsequently post-trained on the curated SynDeC dataset to enhance nuanced argument evaluation. With the enhancement, DebateCV leverages a “star chamber” among LLM agents to accurately verify claims.
@@
Experiments show that DebateCV outperforms existing SOTA methods by 2.6–5.8% in accuracy across diverse evidence quality settings. Case studies illustrate that DebateCV’s adversarial debate mechanism effectively corrects typical single-agent errors, such as claim/evidence misinterpretation, overlooked evidence, and overreliance on speculation. Human expert evaluations further confirm that DebateCV provides higher-quality justifications than previous methods, beyond improvements in accuracy.
@@ 2. Related Work
Automated fact-checking systems (Rothermel et al., 2024; Yoon et al., 2024; Wang et al., 2024; Guo et al., 2022; Leippold et al., 2025; Ullrich et al., 2024) typically operate in two stages: first, they retrieve relevant evidence (evidence retrieval stage); then, they use a single LLM agent to predict a verdict, such as Supported, Refuted, Not Enough Evidence, or Conflicting Evidence/Cherry-picking, along with a justification (Schlichtkrull et al., 2024a; Yoon et al., 2024) (claim verification stage). [^4] Appendix A provides the definition of the four verdict categories.
However, as shown in Sec. 1, existing single-agent systems are vulnerable to multifaceted and even misleading evidence retrieved from the open web. This vulnerability can arise from a fundamental mismatch: LLMs are pretrained to model linguistic patterns by autoregressively predicting the next token, while claim verification demands comprehensive analysis of collected evidence (Mohsin et al., 2025). Ensembling multiple agents, e.g., via majority voting (Choi et al., 2025; Lin et al., 2025) or collaborative frameworks (Li et al., 2023a), also lacks rigorous adversarial scrutiny and may amplify individual errors instead of correcting them. Inspired by real-world fact-checking practices (Graves, 2013), we propose the first debate-driven claim verification framework, DebateCV, to expose and resolve individual assessment errors.
Close to our work, a few recent studies (Wang et al., 2024; Yue et al., 2024) have explored the value of integrating evidence or arguments from opposing perspectives. Wang et al. (2024) retrieve and present separate sets of supporting and opposing evidence to a single LLM for verdict prediction, which focuses on improving evidence retrieval. Instead, we target the subsequent claim verification stage by enabling comprehensive analysis of retrieved evidence. Yue et al. (2024) generate static supporting and refuting arguments independently, but their approach lacks an interactive debate mechanism to surface flaws or resolve contradictions, a mechanism that DebateCV mimics from human fact-checking. Another concurrent work (Han et al., 2025a) investigates evidence-based multi-agent debate for misinformation detection, but it is limited to binary (true/false) classification. This oversimplifies real-world scenarios where claims may lack sufficient evidence (Not Enough Evidence) or involve ambiguity and source disagreement (Conflicting Evidence/Cherry-picking) (Schlichtkrull et al., 2024b). Also, they rely solely on zero-shot LLMs, which, as shown in Sec. 5.3, exhibit conformity bias and frequently default to neutral verdicts in realistic non-binary settings. In contrast, we introduce Debate-SFT, the first post-training framework that substantially improves LLMs’ ability to adjudicate complex debates. We are also the first to conduct a human expert evaluation of justification quality (see Sec. 5.2) to demonstrate the superior transparency of DebateCV framework.
@@
Multi-agent interactions have empowered diverse applications, such as AI societies (Li et al., 2023a) and software engineering (Guo et al., 2024; Hong et al., 2024; Wu et al., 2024), with multi-agent debate-based frameworks often outperforming other forms in complex tasks, such as machine translation (Liang et al., 2024), hallucination mitigation (Du et al., 2024; Hu et al., 2025), and stance detection (Lan et al., 2024). However, prior approaches to misinformation detection (Jeptoo and Sun, 2024; Han et al., 2025b; Liu et al., 2025) mainly debate linguistic patterns of fake news without retrieving external evidence. These patterns are typically platform-specific (Li et al., 2024b) or time-constrained (Yang et al., 2024), limiting their real-world applicability. In contrast, this work targets the claim verification task that analyses consistencies and contradictions between retrieved evidence and claims (Guo et al., 2022), and is the first to propose a debate-driven framework for this task.
Beyond downstream applications, three prior studies have explored leveraging debate mechanisms for post-training LLMs. Subramaniam et al. (2024) employ debates to refine training corpus annotations for general language modelling. Li et al. (2024a) use debates to generate convincing arguments for fine-tuning LLM persuasiveness. Estornell et al. (2024) use debate games between two LLM teams to improve their team-working abilities. Different from these works, we propose Debate-SFT, the first post-training framework specifically designed for enhancing LLMs’ ability to distil multi-round debates into accurate verdicts. We note that, though tailored for claim verification, Debate-SFT, with minimal modifications, can be adapted to other contexts using multi-agent debate.
@@ 3. DebateCV: Debate-driven Claim Verification
Debate-driven claim verification leverages multi-round debates for thorough evidence analysis. We formalize this task as determining the veracity of a textual claim $c$ using an evidence set $\mathcal{E}=\{e_{1},\ldots,e_{k}\}$ by simulating a debate between specialised LLM agents. The outcome is a verdict $\hat{y}$ and a corresponding justification $\hat{j}$, both grounded in the arguments exchanged during the debate. In this section, we present DebateCV and describe how LLM agents are coordinated within the framework.
@@ 3.1. Pre-debate Configuration
Similar to debates conducted by humans, debates conducted by LLM agents must be carefully configured through role assignment. DebateCV employs three distinct LLM agents with complementary responsibilities: two Debaters, namely an Affirmative Debater $D^{+}$ and a Negative Debater $D^{-}$, and one Moderator $M$. Each agent is initialized with tailored system prompts that define their behaviour throughout the debate. Their roles are outlined as follows.
@@ 3.2. In-Debate Process
At the end of each round, the Moderator ($M$) synthesises the key arguments and evidence from both sides into a summary of all exchanged arguments ($S_{t}$). The Moderator then assesses whether the debate has converged, e.g., it deems that $S_{t}$ and $S_{t-1}$ are highly similar, indicating no new insights are being introduced. The debate terminates early if $M$ detects convergence; otherwise, it proceeds until the maximum number of rounds ($T_{\text{max}}$) is reached. Upon conclusion, $M$ determines the final verdict ($\hat{y}$) by weighing the arguments from both sides and provides a justification ($\hat{j}$).
The specific prompts used are provided in Appendix B.
@@ 4. Debate-SFT: Post-training the Moderator
The DebateCV framework, introduced in Sec. 3, provides a debate-driven approach to claim verification. In our preliminary experiments, we find that zero-shot Debater agents consistently produce well-reasoned arguments from their assigned perspectives (see Sec. 5.4 for human evaluation results regarding debate quality). However, the zero-shot Moderator often struggles with complex claims that require multiple debate rounds and tends to exhibit conformity bias (Weng et al., 2025; Zhang et al., 2024). Specifically, it frequently defaults to neutral verdicts such as Not Enough Evidence or Conflicting Evidence/Cherry-picking, even when sufficient evidence exists to support a definitive judgement (see Sec. 5.3 for experimental results).
To enhance the LLM agents’ ability to adjudicate debates, we aim to employ supervised fine-tuning (SFT), a well-established approach for adapting LLMs to downstream tasks (Lai et al., 2025). However, SFT requires annotated debate-driven claim verification datasets, which contain complete multi-round debate transcripts paired with final verdicts and justifications grounded in the debates. Such datasets are both scarce and expensive to curate. To address this challenge, we introduce Debate-SFT, a novel post-training framework that transforms existing non-debate claim verification datasets into synthetic debate-driven equivalents, enabling effective fine-tuning of the Moderator. This section outlines the three stages of Debate-SFT: (1) debate data synthesis, (2) error correction, and (3) post-training.
@@
Given a non-debate claim verification dataset with $n$ claims $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, their ground-truth verdicts $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$, and golden evidence sets $\mathcal{E}=\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{n}\}$ annotated by humans, we leverage the zero-shot DebateCV framework to generate synthetic debate recordings $\mathcal{D}=\{d_{1},d_{2},\dots,d_{n}\}$. Each recording $d_{i}$ is a multi-round transcript of the arguments and summaries produced during the debate over claim $c_{i}$: $d_{i}=\bigl\{A^{+}_{1},A^{-}_{1},\dots,A^{+}_{t_{i}},A^{-}_{t_{i}}\bigr\},$ where $A^{+}_{t_{i}}$ and $A^{-}_{t_{i}}$ are the affirmative and negative arguments in round $t_{i}$}. [^5] Our human evaluation in Sec. 5.4 shows that the generated debates feature well-reasoned, evidence-based arguments and consistently uphold opposing stances. During this process, the zero-shot Moderator predicts verdicts $\mathcal{\hat{Y}}=\{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{n}\}$ with justifications $\mathcal{\hat{J}}=\{\hat{j}_{1},\hat{j}_{2},\dots,\hat{j}_{n}\}$ based on $\mathcal{D}$.
@@
Since the zero-shot Moderator may issue erroneous verdicts even when the debaters’ arguments are established in gold-standard evidence, we use human-rated ground-truth verdicts to ensure data validity. Specifically, when the predicted verdict $\hat{y}_{i}$ differs from the ground-truth $y_{i}$, we introduce an LLM-based Corrector to generate a justification that aligns with the human annotation. For each claim $c_{i}$ where $\hat{y}_{i}\neq y_{i}$, the Corrector takes the full debate recording $d_{i}$ and the human-rated ground-truth verdict $y_{i}$ as input to generate a corrected justification $j^{\prime}_{i}$ that explains how $y_{i}$ can be derived from the arguments and evidence in $d_{i}$. [^6] Our human evaluation in Sec. 5.4 shows that these Corrector-generated justifications effectively reflect the debated content and align with human judgements on the verdict.
@@
where $\mathcal{L}_{\text{SFT}}$ is the standard SFT loss (Ouyang et al., 2022) and $\theta$ represents the zero-shot Moderator parameters. This three-stage Debate-SFT framework results in an enhanced Moderator that can better evaluate debate dynamics and render more accurate judgements.
@@
To instantiate the Debate-SFT framework, we apply it to the training split of AVeriTeC (Schlichtkrull et al., 2024b) (3,068 claims) to create the first synthetic dataset designed for debate-driven claim verification, SynDeC. [^7] Our method can be applied to any non-debate claim verification dataset with human-rated verdicts and evidence. Specifically, we use the zero-shot DebateCV framework to synthesise full debate recordings $d_{i}$ for all 3,068 claims. Among these, 936 claims are incorrectly predicted and undergo error correction, where a Corrector based on GPT-4o generates a corrected justification $j^{\prime}_{i}$ aligned with the ground-truth verdict $y_{i}$. Each sample in the SynDeC dataset contains the following elements: (i) claim $c_{i}$, (ii) evidence set $\mathcal{E}_{i}$, (iii) debate recording $d_{i}$, (iv) ground-truth verdict $y_{i}$, (v) predicted verdict $\hat{y}_{i}$, (vi) predicted justification $\hat{j}_{i}$, and, when $\hat{y}_{i}\neq y_{i}$, (vii) the corrected justification $j^{\prime}_{i}$.